# Nash CoT: Multi-Path Inference with Preference Equilibrium

### v1

@@ 3.1 Mini-batch inference with Preference Equilibrium
Subsequently, based on the concept of Preference Equilibrium, we conceptualize a Mini-batch inference (shown in Figure 1) as a bi-player gaming system. This approach aims to achieve better inference compared to direct inference, while still preserving some of the inherent randomness of standard inference methods. To begin with proposing this system, we first define $x^{t}$ as the template of zero-shot CoT, $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ (we provided several cases in Player Templates) as the candidates template for template guided generation. Meanwhile, in this system, the template can to be chosen by a reward model $r_{\theta}$.
[…]
4 Nash Chain of Thought (Nash CoT)
@@
Nash CoT can be seen as an extension of Mini-batch inference with Preference Equilibrium, implementing multiple Mini-batch inferences to enhance performance. This approach is influenced by experimental results from self-consistency, which suggest that increasing the number of paths can improve inference accuracy. Meanwhile, the reasoning process for each question is divided into two stages: Answer Gathering and Answer Filtering.
@@
When generating candidate answers, the process predominantly involves two types of loops: Mini-batch Loops ($\rm n_{\rm mini}$): In Chapter 3.1, we discussed the implementation of mini-batch inference with Preference Equilibrium. As shown in Algorithm 1, this process involves searching for template-guided generations within two rounds of generation ($[y_{1},y_{2}]$). We refer to the times of these two predictions as the $\rm n_{\rm mini}$. Moreover, to mitigate the impact brought from low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times. This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path in self-consistency. After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches equilibrium most frequently (shown in Algorithm 2), as the preferred answer.
@@
Mistral-Instruct (7B) zero-shot 15.3 $\pm$ 0.8 12.0 $\pm$ 2.8 3.3 $\pm$ 1.3 2.7 $\pm$ 2.0 20.8 $\pm$ 1.5 7.7 $\pm$ 2.0 10.3$\pm$ 6.5
@@
GLM4-chat (9B) zero-shot 1.1 $\pm$ 1.5 1.1 $\pm$ 1.5 12.6 $\pm$ 3.9 12.0 $\pm$ 2.0 22.4 $\pm$ 4.1 4.4 $\pm$ 2.8 8.9$\pm$7.6
zero-shot CoT 90.7 $\pm$ 1.5 90.7 $\pm$ 1.5 98.4 $\pm$ 1.3 80.9 $\pm$ 2.8 20.8 $\pm$ 3.1 86.9 $\pm$ 3.5 78.1$\pm$26.1
self-consistency (20 Paths) 92.3 $\pm$ 2.0 90.2 $\pm$ 2.3 98.4 $\pm$ 2.3 89.3 $\pm$ 0.2 20.8 $\pm$ 3.1 91.5 $\pm$ 1.1 80.4$\pm$26.8
Nash CoT (10 Paths) 91.3 $\pm$ 0.8 90.2 $\pm$ 2.7 96.7 $\pm$3.3 80.3$\pm$ 1.3 20.8$\pm$3.1 88.0 $\pm$2.0 77.9$\pm$26.1
Table 1: Experimental results on arithmetic reasoning benchmarks. We test Zero-Shot CoT and Nash CoT with the core LLM includes Mistral-Instruct (7B) and GLM4-chat (9B) on mathematical benchmarks including AddSub, MultiArith, SingleEQ, SVAMP, GSM8K, and AQuA. Nash CoT performs the best.
Answer Filtering.
In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction satisfy Preference equilibrium. Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency. Subsequently, we return the most frequent case. Otherwise, if is no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adopt the strategy of self-consistency by selecting the most frequent prediction among all generated answers.
@@
Subsquently, we propose Nash CoT, which iterates through Algorithm 1 and Algorithm 2 to perform inference on all sampled questions, where $\tau$ represents the candidate answers from the Answer Gathering stage.
@@
Practical Implementation of reward model $r_{\theta}$.
In the process of practical implementation, we do not explicitly train a reward model $r_{\theta}$ to confine the player template $x^{c}$ (we have provided cases in Player Templates) using Equation 1. Instead, we directly use the preference template (shown in Preference Template) to guide the LLM in determining the most suitable player template for a given question. For example, when presented with a coin flip question as shown in Figure 4, we fill the Preference Template with given question and player templates. This filled template is then input into the LLM to provide the id of the most suitable player template from the available options. In particular, we believe it’s effective, this is because most of baselines we selected have been turned to reflect human preference, thus we believe the selected LLM can be directly utilized as the preference model to point out the most preferred option among candidate options.
@@ 5 Experiments
Nash CoT (10 Paths) 29.0$\pm$ 5.4 0.5 $\pm$ 0.8 44.8 $\pm$ 2.0 41.1 $\pm$ 1.2 28.9$\pm$ 17
GLM4-chat (9B) zero-shot 27.3 $\pm$ 6.3 0.0 $\pm$ 0.0 38.8 $\pm$ 0.8 16.4 $\pm$ 2.3 20.7$\pm$14.3
zero-shot CoT 87.4 $\pm$ 0.8 0.0 $\pm$ 0.0 37.7 $\pm$ 2.3 16.4 $\pm$ 4.8 35.4$\pm$32.9
self-consistency (20 Paths) 98.9$\pm$1.5 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 38.3$\pm$37.5
Nash CoT (10 Paths) 93.4$\pm$2.7 0.0 $\pm$ 0.0 37.7$\pm$2.3 16.4$\pm$4.8 36.9$\pm$35.3
Table 2: Experimental results on symbolic inference benchmarks. We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Symbolic QA benchmarks includes Coin-Flipping, Last Letters and Object Tracking. Among these baselines, Nash CoT performs the best.
@@
GLM4-chat (9B) zero-shot 56.8 $\pm$ 4.7 17.5$\pm$ 2.0 37.2$\pm$19.7
zero-shot CoT 63.9$\pm$2.3 18.0$\pm$2.3 41.0$\pm$ 22.9
self-consistency (20 Paths) 69.9 $\pm$3.3 18.0$\pm$2.3 44.0$\pm$26.0
Nash CoT (10 Paths) 66.7 $\pm$ 0.8 18.0 $\pm$ 2.3 42.4$\pm$24.4
@@
The goal of our experiment is to 1) demonstrate the performance advantage and effectiveness of Nash CoT. 2) shocase whether Nash CoT help reduce the overall inference time. In the following sections, we first introduce our experimental setup and then present the experimental results and analysis.
@@
Our majority benchmarks are composed of three different kinds of inference tasks. 1) arithmetic reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021). 2) symbolic reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date. 3) commonsense question answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Appendix A.
@@
To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-7B (Instruct) (Jiang et al., 2023), GLM4-9B-Chat Zeng et al. (2022); Du et al. (2022). In particular, all of these selected LLMs are turned via RL with human feedback (RLHF), and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al. (2022).
@@
The preliminary baselines we utilized include zero-shot, zero-shot CoT (Wei et al., 2023), andself-consistency (Wang et al., 2023). We test these approach with freezed LLMs.
@@
Our evaluation on all selected tasks utilizes the same experimental settings bellow:
@@
  • self-consistency. We follow Wang et al. (2023) to evaluate the performance ofself-consistency with selected LLMs, utilizing the zero-shot CoT template. Additionally, we set the number of inference paths to 20.
  • Nash CoT. We set up $n_{\rm outer}$ as 3 and $n_{\rm mini}$ as 2, resulting in a total of $n_{\rm outer}\times(n_{\rm mini}+1)+1=10$ paths. Additionally, we have provided the player templates $x^{t}$ in Table 1 and the Appendix, meanwhile we utilizing the same CoT template $x^{c}$ as in zero-shot CoT.
Additionally, all evaluations are conducted on the inference of 60 random sampled questions multi times. And we have provided the mean and standard error in all tables.
Figure 3: We used GLM4-chat (9B) on the same type of GPU (A-100) to evaluate Nash CoT and self-consistency across selected tasks. Nash CoT, employing a total of 10 paths, requires nearly half the time of self-consistency, which has 20 paths in total.
@@ 5.1 Experimental Results
The majority experimental results are demonstrated in table 12 and 3. Nash CoT can improve Mistral-Instruct (7B) on almost all selected inference tasks, while showcasing similar performance to self-consistency with twice inference paths on GLM4-chat (9B). In particular, we have provided the total paths of Nash CoT that it only require the half of self-consistency, thus our claim in section 3 can be validated.
When focusing on Mistral-Instruct (7B), Nash CoT has better performance on arithmetic and symbol inference tasks, showcasing its superiority performance on logic/math inference tasks. However, Nash CoT does not showcase improved performance in commonsense question answering tasks. We argue that this is because commonsense question answering tasks are more diverse, and the player template can’t cover all topics. Therefore, the player template limits Nash CoT on commonsense question answering tasks. Importantly, we limit Nash CoT’s performance by utilizing only total 10 paths for inference in this section. However, additional experimental results in the ablation section show that Nash CoT outperforms self-consistency via increasing the inference loops, thus Nash CoT can outperform self-consistency.
@@ 6 Ablation Study
In order to further validate the effectiveness of Nash CoT, we conducted extensive ablations to answer the following questions: 1) What will happen when the number of inference paths for Nash CoT is further increased? Will Nash CoT eventually surpass self-consistency, and what is the relationship between the number of loops and performance? 2) Does the template really improve the accuracy of path predictions, and what impact does it have on experimental performance?
@@
To address question 1), we selected Mistral-Instruct (7B) and conducted evaluation on three different reasoning tasks, adjusting the $N_{\rm mini}$ and $N_{\rm outer}$. As shown in Figure 4, as the number of loops increases, Nash CoT has a high probability of significantly outperforming self-consistency with fewer paths. However, different from self-consistency, the experimental results of Nash CoT do not show a monotonic (linear) relationship with the total number of total paths. This indicates that there is a significant difference between Nash CoT and self-consistency. Unlike Nash CoT, the experimental results of self-consistency show a clear improvement in performance as the number of paths increases.
@@
To illustrate the impact of the template, we removed the mathematical templates from the Player Templates and then evaluated Nash CoT on selected Arabic reasoning. Results are shown in Table 4, showing an approximately 9.2% decrement in GSM8K and 6.2% decrement in SVAMP. Therefore, the performance of Nash CoT is impacted by the Player Template.

### v2

@@ 3.1 Mini-batch inference with Preference Equilibrium
Subsequently, based on Preference Equilibrium, we conceptualize a mini-batch inference (shown in Step 2 of Figure 1) as a bi-player gaming system. This approach aims to maintain the performance benefits of multi-path inference while retaining some of the inherent randomness (diversity) of standard inference methods. Before introducing mini-batch inference, we first define $x^{t}$ as the template of zero-shot CoT. Meanwhile, we leverage LLMs to select a player template from the candidate set $\{x^{c}_{0},x^{c}_{1},\cdots,x^{c}_{n}\}$ for role-immersed generation, guided by Equation 1.
@@
[…]
Based on the mini-batch inference, we further introduce Nash CoT in the next chapter. (In particular, in the current framework, successful inference cannot always be guaranteed. For instance, $y^{*}$ may not always in $[y_{1},y_{2}]$. We will address this problem in the following chapters.)
4 Nash Chain of Thought (Nash CoT)
Nash CoT can be regarded as an extension of mini-batch inference with Preference Equilibrium, implementing multiple mini-batch inferences to further enhance performance. This approach is inspired by self-consistency experiments, which suggest that increasing the number of paths enhances LLM’s inference accuracy. Specifically, Nash CoT’s reasoning process for each question can be divided into two stages: Answer Gathering and Answer Filtering.
@@ 4 Nash Chain of Thought (Nash CoT)
When generating candidate answers, the process predominantly involves two types of loops: mini-batch Loops ($\rm n_{\rm mini}$): As shown in Algorithm 1, this process involves searching for role-immersed generations within two rounds of generation $[y_{1},y_{2}]$. We denote the times of these two predictions as the $\rm n_{\rm mini}$. Moreover, to mitigate the impact brought by low-frequency predictions, we introduce iterating $\rm n_{\rm mini}$ multiple times. This leads us to another type of loop: Outer Loops ($\rm n_{\rm outer}$): This loop resembles the concept of multi-path inference that iterates $\rm n_{\rm mini}$ multiple times. After completing loop $\rm n_{\rm outer}$, we filter the generated answers and retain the answer that reaches Preference Equilibrium most frequently (shown in Algorithm 2), as the predicted answer.
@@
Mistral-Instruct (7B) zero-shot 15.3$\pm$0.8 12.0$\pm$2.8 3.3$\pm$1.3 2.7$\pm$2.0 20.8$\pm$1.5 7.7$\pm$2.0 10.3$\pm$7
@@
GLM4-chat (9B) zero-shot 1.1$\pm$1.5 1.1$\pm$1.5 12.6$\pm$3.9 12.0$\pm$2.0 22.4$\pm$4.1 4.4$\pm$2.8 8.9$\pm$8
zero-shot CoT 90.7$\pm$1.5 90.7$\pm$1.5 98.4$\pm$1.3 80.9$\pm$2.8 20.8$\pm$ 3.1 86.9$\pm$3.5 78.1$\pm$26
self-consistency (20 Paths) 92.3$\pm$2.0 90.2$\pm$2.3 98.4$\pm$2.3 89.3$\pm$0.2 20.8$\pm$3.1 91.5$\pm$1.1 80.4$\pm$27
Nash CoT (10 Paths) 91.3$\pm$0.8 90.2$\pm$2.7 96.7$\pm$3.3 80.3$\pm$1.3 20.8$\pm$3.1 88.0$\pm$2.0 77.9$\pm$26
Table 1: Experimental results on Arabic Reasoning benchmarks. We test Zero-Shot CoT and Nash CoT with the core LLM includes Mistral-Instruct (7B) and GLM4-chat (9B) on mathematical benchmarks including AddSub, MultiArith, SingleEQ, SVAMP, GSM8K, and AQuA. Nash CoT performs the best.
@@
Answer Filtering.
In terms of answer filtering, as shown in Algorithm 2 we first count the most frequent prediction to satisfy Preference equilibrium. Specifically, we count all $y^{*}$ satisfy $y^{*}\in[y_{1},y_{2}]$ and compute their frequency. Subsequently, we return the most frequent answer. Otherwise, if no cases satisfy $y^{*}\in[y_{1},y_{2}]$, we adapt answer filtering by selecting the most frequent prediction among all generated answers.
@@
Practical Implementation of Preference model $\mathcal{P}$.
In the process of practical implementation, we do not explicitly train a reward model $r_{\theta}$ to confine the player template $x^{c}$ utilizing Equation 1. Instead, we directly utilize the template (shown in Preference Template) to guide the LLM in determining the most suitable player template for a given question. For example, when presented with a coin flip question as shown in Figure 4, we fill the Preference Template with the given question and Player Templates. This filled template is then input into the LLM to provide the id of the most suitable player template from the available options. In particular, we consider utilizing LLM to search templates to be effective, this is because most of the baselines we selected have been turned to reflect human preference, thus LLM can be directly utilized as the preference model to point out the most preferred option among candidate options (we provide cases in section A). Subsequently, we propose Nash CoT, which iterates through Algorithm 1 and Algorithm 2 to perform inference on all sampled questions, where $\tau$ represents the candidate answers from the Answer Gathering stage.
@@ 5 Experiments
Nash CoT (10 Paths) 29.0$\pm$5.4 0.5$\pm$0.08 44.8$\pm$2.0 41.1$\pm$1.2 28.9$\pm$17
GLM4-chat (9B) zero-shot 27.3$\pm$6.3 0.0$\pm$0.0 38.8$\pm$0.8 16.4$\pm$2.3 20.7$\pm$14
zero-shot CoT 87.4$\pm$0.8 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 35.4$\pm$33
self-consistency (20 Paths) 98.9$\pm$1.5 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 38.3$\pm$38
Nash CoT (10 Paths) 93.4$\pm$2.7 0.0$\pm$0.0 37.7$\pm$2.3 16.4$\pm$4.8 36.9$\pm$35
Table 2: Experimental results on symbolic inference benchmarks. We test Zero-Shot CoT and Nash CoT with Mistral-Instruct (7B) and GLM4-chat (9B) on Symbolic QA benchmarks including Coin-Flipping, Last Letters, and Object Tracking. Among these baselines, Nash CoT performs the best.
@@
GLM4-chat (9B) zero-shot 56.8$\pm$4.7 17.5$\pm$2.0 37.2$\pm$20
zero-shot CoT 63.9$\pm$2.3 18.0$\pm$2.3 41.0$\pm$23
self-consistency (20 Paths) 69.9$\pm$3.3 18.0$\pm$2.3 44.0$\pm$26
Nash CoT (10 Paths) 66.7$\pm$0.8 18.0$\pm$2.3 42.4$\pm$24
@@
The goal of our experiment is to 1) showcase the performance advantage and effectiveness of Nash CoT. 2) showcase whether Nash CoT helps reduce inference paths and inference time.
@@
Our majority benchmarks are composed of three different kinds of inference tasks. 1) Arabic Reasoning: SingleEq (Koncel-Kedziorski et al., 2015), AddSub (Hosseini et al., 2014), MultiArith (Roy and Roth, 2016), GSM8K (Cobbe et al., 2021), AQUA (Ling et al., 2017), and SVAMP (Patel et al., 2021). 2) Symbolic Reasoning: Last Letters, Coin Flip (Wei et al., 2023), and Object Tracking, Bigbench Date. 3) Commonsense Question Answering: CommonsenseQA Talmor et al. (2019) and StrategyQA Geva et al. (2021). For more details about the dataset please refer to Appendix A.
@@
To validate that Nash CoT is a general CoT method, we selected different large models as test models, including Mistral-Instruct (7B) (Jiang et al., 2023), Chat-GLM4 (9B) Zeng et al. (2022); Du et al. (2022). In particular, all of these selected LLMs are turned via RLHF, and the difference between LLM turned with RLHF and the original foundation models have been detailed by Ouyang et al..
@@
The preliminary baselines we utilized include zero-shot, zero-shot CoT (Wei et al., 2023), and self-consistency (Wang et al., 2023). We test these approaches with frozen LLMs.
@@
Our evaluation of all selected tasks utilizes the same experimental settings below:
@@
  • self-consistency. We follow Wang et al. to evaluate the performance of self-consistency with selected LLMs, utilizing the zero-shot CoT template. Additionally, we set the number of inference paths to 20.
  • Nash CoT. We set up $n_{\rm outer}$ as 3 and $n_{\rm mini}$ as 2, resulting in a total of $n_{\rm outer}\times(n_{\rm mini}+1)+1=10$ paths. Additionally, we have provided the Player Templates $x^{t}$ in Table 1 and the Appendix, meanwhile, we utilizing the same CoT template $x^{c}$ as in zero-shot CoT.
Specifically, in Table 12 and 3, we set the total number of reasoning paths for Nash CoT to 10 and for self-consistency to 20. The primary goal is to demonstrate that Nash CoT can achieve the same or even better results with only half the number of reasoning paths compared to self-consistency.
Additionally, all evaluations are conducted on the inference of 60 random sampled questions per seed. We have provided the mean and standard error in the majority tables.
Figure 3: We used GLM4-chat (9B) on the same type of GPU (A100) to evaluate Nash CoT and self-consistency across selected tasks (60 questions per task). Nash CoT, employing a total of 10 paths, requires nearly half the time of self-consistency, which has 20 paths in total.
@@ 5.1 Experimental Results
Models Settings Cnt and Prob Precalculus Inter-Alg Number Theory Avg.
Mistral-Instruct (7B) zero-shot-cot 40.9 16.4 16.4 24.6 24.6
Nash CoT 64.8 38.5 36.9 27.9 42.0
self-consistency 72.1 34.4 34.4 21.3 40.6
GLM4 (large) zero-shot-cot 83.6 90.2 85.2 91.9 87.7
Nash CoT 95.1 90.2 98.3 98.3 95.5
self-consistency 93.3 90.2 95.1 100.0 94.7
Table 4: Equal Path Evaluation. This experiment is conducted to verify the performance of Nash CoT when the number of inference paths is set equal to self-consistency. In particular, Cnt and Prob denote Counting and Probability, while Inter-Alg denotes Intermediate Algebra.
@@
The majority experimental results are demonstrated in Table 12 and 3. Nash CoT can improve Mistral-Instruct (7B) on almost all selected inference tasks while showcasing similar performance to self-consistency with twice inference paths on GLM4-chat (9B). In particular, we have provided the total paths of Nash CoT that it only require half of self-consistency, thus our claim in section 3 can be validated. When focusing on Mistral-Instruct (7B), Nash CoT has better performance on Arabic and Symbolic inference tasks, showcasing its superior performance on logic/math inference tasks. However, Nash CoT does not showcase improved performance in Commonsense Question Answering tasks. We argue that this is because Commonsense Question Answering tasks are more diverse, and the player template can’t cover all topics. Therefore, the predefined player template limits Nash CoT on Commonsense Question Answering tasks. Importantly, we evaluate Nash CoT by utilizing only a total of 10 paths for inference in this section. However, additional experimental results in the ablation section show that Nash CoT outperforms self-consistency by increasing the inference loops. Meanwhile, we also evaluate by setting equal paths on much more challenging MATH Hendrycks et al. (2021), showcasing that Nash CoT can outperform self-consistency.
Comparing Nash CoT with self-consistency with equal paths.
To further evaluate the performance of Nash CoT, we utilize Mistral-Instruct (7B) while selecting one of state-of-the-art (SOTA) LLMs named GLM4 (large) (GLM et al., 2024), for comparison using an equal number of reasoning paths (10 paths). As shown in Table 4, Nash CoT outperforms self-consistency, which demonstrates the superiority of Nash CoT compared to self-consistency when the same number of path settings are set.
@@ 6 Ablation Study
We further conduct extensive ablations to answer the following questions: 1) Is there a correlation between the performance of Nash CoT and the setting of the number of inference paths and can the performance of Nash CoT be improved to surpass self-consistency by adjusting the number of inference paths? 2) Does the template improve the accuracy of path predictions, and what impact does it have on experimental performance? 3) Is the Nash Equilibrium the reason for the performance improvement observed in Nash CoT?
@@
To address question 1), we selected Mistral-Instruct (7B) and evaluated three different reasoning tasks, adjusting the $N_{\rm mini}$ and $N_{\rm outer}$. As shown in Figure 4, as the number of loops increases, Nash CoT has a high possibility of outperforming self-consistency with fewer paths. However, different from self-consistency, the experimental results of Nash CoT do not show a monotonic (linear) relationship with the total number of paths. This indicates that there is a significant difference between Nash CoT and self-consistency. Unlike Nash CoT, the experimental results of self-consistency show a clear improvement in performance as the number of paths increases.
@@
To illustrate the impact of the template, we removed the mathematical templates from the Player Templates and then evaluated Nash CoT on selected Arabic reasoning. Results are shown in Table 5, showing an approximately 9.2% decrement in GSM8K and 6.2% decrement in SVAMP. Therefore, the performance of Nash CoT is impacted by the Player Templates.
@@
As shown in Table 6, we used three role templates with varying degrees of relevance to mathematical problems: mathematician, student, and poker player, to test the performance of Mistral-Instruct (7B) on GSM8K. Role templates with higher relevance to mathematical problems can bring more improvements to Mistral, further confirming the rationality of enhancing reasoning performance by substituting LLM into roles using templates.
mathematician student poker player
57.37 52.45 47.54
Table 6: Results on GSM8K. The templates for these roles are shown in Table 9 on the Appendix.
Effectiveness of Nash Equilibrium.
While Nash CoT has demonstrated effectiveness across various LLMs and benchmarks, however, the extent to which Nash CoT’s performance is driven by Nash Equilibrium remains to be determined. To further investigate, we compare the self-consistency integrated template (denoted as self-con.+temp (10 paths)) with Nash CoT (10 paths) using Mistral-Instruct (7B). As shown in Table 7, Nash CoT consistently outperforms self-con.+temp when applying the same mathematical templates across selected tasks.
Settings AQuA SVAMP GSM8K AddSub Avg.
self-con.+temp 31.7 71.6 53.6 85.8 60.6
Nash CoT 39.9 77.0 55.7 86.3 64.7
Table 7: Ablation of Nash Equilibrium. This experiment is conducted to verify the necessity of Nash Equilibrium in the performance of Nash CoT.